How to solve 774+447+113

Found 2 solutions by ikleyn, MathLover1:
Answer by ikleyn(52778)

.

On adding three-digit numbers see,  read and learn from this web-site

http://www.numbernut.com/arithmetic/add-3digit.html



Answer by MathLover1(20849)


you just add them

if you are given specific base, for example base 8, then you will have
774
447
113
-------- Add
4+7+3 = 6 carry 1
7+4+1+1 = 5 carry 1
7+4+1+1 = 5 carry 1
--> 1556